# Hydrostatic mass of galaxy clusters within some theories of gravity

### v4

@@ 1 Introduction
Galaxy clusters (GCs) are the largest virialized astronomical structures in the universe [1]. They are generally defined as self-gravitating systems of galaxies that may be identified in observations as the systems exhibiting one or more of the following features: a significantly higher concentration of galaxies relative to the average distribution of the large-scale structure of the universe; massive concentration(s) of X-ray emitting gas, often associated with the Sunyaev-Zel’dovich (SZ) effect on the Cosmic Microwave Background Radiation; and a non-random gravitational lensing shear map of background galaxies [2]. As a relatively large structure, GCs are magnificent laboratories for constraining fundamental physics, such as bounding the neutrino mass, graviton mass, or limits on primordial non-Gaussianity. GCs also have an important role in cosmology due to the many discoveries based on optical, X-ray, and microwave surveys. [3].
@@ 2.1 Hydrostatic mass of GCs within EiBI theory
$M_{EiBI}(r)=-\left[3,7\times 10^{-4}\>T\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)+\frac{r\kappa\rho}{4M_{\odot}}\left(\frac{d\ln\rho}{d\ln r}\right)\right]\>\mathscr{M}_{\odot}.$ (6)
@@ 2.2 Hydrostatic mass of GCs within BHG
$\frac{M_{BHG}(r)}{\mathscr{M}_{\odot}}=-\frac{rk_{b}T}{\mu m_{p}G}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{M_{\odot}}\left(2+\frac{d\ln\rho}{d\ln r}\right).$ (12)
By substituting the values of the constants, we finally find the mass of GCs within BHG in the unit of $M_{\odot}$, i.e.
$M_{BHG}=\left[{-3,7\times 10^{-4}}\>T\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{M_{\odot}}\left(2+\frac{d\ln\rho}{d\ln r}\right)\right]\mathscr{M}_{\odot}.$ (13)
@@ 3 Numerical results
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$ m2 for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [19], [43], and [36]. In Ref. [19], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [43], $\kappa=0.2$ m2 could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we realize that the objects used are white dwarf in Ref. [19], polytropic star in Ref. [43], and brown dwarf in Ref. [36]. Compared to GCs, those objects have different structures and EOS. So it is clear that GCs have their own values of the free parameters (i.e. $\beta_{0}$, $\kappa$, and $\Upsilon$) to conduct some impact on the mass. However, although $\beta_{0}$ is a free parameter, setting its value to produce a significant impact observable on astronomical scales is not appropriate, as $\beta_{0}$ originates from quantum corrections.
@@
It is important to note that for both EiBI theory and BHG when we set $\kappa$ and $\Upsilon$ as positive numbers, the mass of GCs decreases. As we magnify both $\kappa$ and $\Upsilon$, the hydrostatic mass of GCs increases. Based on this mathematical fact, a trivial thought might arise that we should magnify the values of $\kappa$ and $\Upsilon$ to obtain better linear fits. But in fact, it will not happen, since if we set $\kappa$ and $\Upsilon$ exceed the values used in the linear fits shown by Fig. 2(a) and Fig. 2(b), some masses of GCs reach negative values, which is not physical.

### v5

@@ 1 Introduction
Galaxy clusters (GCs) are the largest virialized astronomical structures in the universe [1]. They are generally defined as self-gravitating systems of galaxies that may be identified in observations as the systems exhibiting one or more of the following features: a significantly higher concentration of galaxies relative to the average distribution of the large-scale structure of the universe; massive concentration(s) of X-ray emitting gas, often associated with the Sunyaev-Zel’dovich (SZ) effect on the Cosmic Microwave Background Radiation; and a non-random gravitational lensing shear map of background galaxies [2]. As a relatively large structure, GCs are magnificent laboratories for constraining fundamental physics, such as bounding the neutrino mass, graviton mass, or limits on primordial non-Gaussianity. GCs also have an important role in cosmology due to the many discoveries based on optical, X-ray, and microwave surveys [3].
@@ 2.1 Hydrostatic mass of GCs within EiBI theory
$M_{EiBI}(r)=-\left[3,7\times 10^{-4}\>T\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)+\frac{r\kappa\rho}{4{\mathscr{M}_{\odot}}}\left(\frac{d\ln\rho}{d\ln r}\right)\right]\>\mathscr{M}_{\odot}.$ (6)
@@ 2.2 Hydrostatic mass of GCs within BHG
$\frac{M_{BHG}(r)}{\mathscr{M}_{\odot}}=-\frac{rk_{b}T}{\mu m_{p}G}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{{\mathscr{M}_{\odot}}}\left(2+\frac{d\ln\rho}{d\ln r}\right).$ (12)
By substituting the values of the constants, we finally find the mass of GCs within BHG in the unit of ${\mathscr{M}_{\odot}}$, i.e.
$M_{BHG}=\left[{-3,7\times 10^{-4}}\>T\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{{\mathscr{M}_{\odot}}}\left(2+\frac{d\ln\rho}{d\ln r}\right)\right]\mathscr{M}_{\odot}.$ (13)
@@ 3 Numerical results
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$ m2 for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [19], [43], and [36]. In Ref. [19], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [43], $\kappa=0.2\times 10^{9}$ m2 could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we realize that the objects used are white dwarf in Ref. [19], polytropic star in Ref. [43], and brown dwarf in Ref. [36]. Compared to GCs, those objects have different structures and EOS. So it is clear that GCs have their own values of the free parameters (i.e. $\beta_{0}$, $\kappa$, and $\Upsilon$) to conduct some impact on the mass. However, although $\beta_{0}$ is a free parameter, setting its value to produce a significant impact observable on astronomical scales is not appropriate, as $\beta_{0}$ originates from quantum corrections.
It is important to emphasize that the choice of the sign of free parameters in Table 1 is not strictly governed by any specific rule. The primary purpose of Table 1 is to identify the parameter values, particularly those farther from zero, at which their impact on the mass within modified gravity theories becomes more pronounced. Moreover, since $\kappa$, $\Upsilon$, and $\beta_{0}$ are free parameters, it is evident from Refs. [19, 36, 43] that these parameters can take either positive or negative values.
@@
It is important to note that for both EiBI theory and BHG when we set $\kappa$ and $\Upsilon$ as positive numbers, the mass of GCs decreases. As both $\kappa$ and $\Upsilon$ are magnified with negative signs, the hydrostatic mass of GCs increases. The parameter signs used in Fig. 2(a) and Fig. 2(b) are intended to reduce the mass, bringing it into alignment with the baryonic mass order, specifically around $10^{13}M_{\odot}$ (see Ref. [41]). As shown in all presented figures, most $M_{bar}$ values fall within this range. These values are notably lower than the mass orders predicted by Newtonian gravity and modified gravity theories used in this work, i.e. EiBI theory and BHG, which are typically around $10^{14}M_{\odot}$. The parameter $\kappa$ is consistently chosen to be positive, as this positive value reduces the GCs’ masses compared to those obtained using Newtonian gravity.
Similarly, the purpose of selecting a positive value for $\Upsilon$ in the BHG theory, which differs from the notation in the table (as the table uses a negative value), is to decrease the mass values, thereby achieving the baryonic mass order. The value $\Upsilon=1.65\times 10^{71}$ sufficiently reduces the masses of some GCs to the baryonic mass order. However, the resulting linear fit slope remains below the slope of Newtonian linear fit. This can be understood by examining Fig. 2(b). While some of the GCs’ masses in the BHG framework fall within the order of the baryonic mass, their distribution does not display a linear pattern as seen in the cases of $M_{Newtonian}$ shown in Fig. 1 and $M_{EiBI}$ shown in Fig. 2(a). Consequently, the slope of the linear fit for the ratio $M_{BHG}/M_{bar}$ has the lowest value among the cases considered. These findings indicate that even the BHG theory is less effective than Newtonian gravity in addressing the mass discrepancy issue in GCs.
Based on the mathematical fact that $M_{EiBI}$ and $M_{BHG}$ decrease when $\kappa$ and $\Upsilon$ are assigned negative values, a trivial thought might arise suggesting that magnifying the negative values of $\kappa$ and $\Upsilon$ could yield better linear fits. But in fact, it will not happen, since if we set $\kappa$ and $\Upsilon$ exceed the values used in the linear fits shown by Fig. 2(a) and Fig. 2(b), some masses of GCs reach negative values, which is not physical.